You bought a stock three months ago for $43.48 per share. The stock paid no dividends. The current share price is $47.09

  

What is the APR of your investment? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places (e.g., 32.16).)

  

  APR %

  

What is the EAR of your investment? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places (e.g., 32.16).)

  

  EAR %
Answer #1

Answer

Answer 1

Figures in $

Particulars

Amount

Current share price

a

47.09

Purchase price

b

43.48

3 month gain      (a-b)

c

3.61

12 month gain     (c*4)

d

14.44

Annual percent rate (d/b)*100

33.21%

Answer 2

Effective Annual Rate = (1 + Annual percent rate/number of compounding period)n – 1

                                       = (1 + 0.3321/4)4 – 1

                                       = (1 + 0.083025)4 – 1

                                        =(1.083025)4 – 1

                                        = 1.3758 -1

                                         = 0.3758

Effective Annual Rate = 37.58%